# Robust Tail Index Estimation under Random Censoring via Minimum Density Power Divergence

### v5

@@ 4.4. Simulation scenarios and settings
Although infinitely many combinations of target, censoring, and contamination distributions could be considered, we restrict our attention to two representative scenarios that are commonly used in extreme value analysis and that allow for a clear and informative assessment of both efficiency and robustness.
The following scenarios are implemented:
@@
  • S1: Burr/Burr/Burr: the target variable $X$ follows Burr distribution, the censoring variable is also Burr-distributed, and contamination is introduced via a Burr distribution with parameters $\eta_{c}=0.25$ and $\gamma_{c}=\left\{0.6,0.8\right\}.$
@@
  • S2: Fréchet/Fréchet/Fréchet: the target variable $X$ follows a Fréchet distribution, the censoring mechanism is Fréchet-distributed, and contamination is generated from a Fréchet distribution with tail index $\gamma_{c}=\left\{0.6,0.8\right\}.$
The MDPD estimator $\widehat{\gamma}_{1,\alpha}$ is computed for $\alpha\in\{0.1,0.3,0.5\}.$ For comparison, we include the following benchmark estimators:
  • Adapted Hill-type $\widehat{\gamma}_{1}^{(EFG)},$
  • Kaplan–Meier-based $\widehat{\gamma}_{1}^{(W)},$
  • Nelson–Aalen-based $\widehat{\gamma}_{1}^{(MNS)}$ (for $\alpha=0).$
This selection covers both classical and censoring-adapted tail index estimators that are widely used in the literature.
4.5. Selection of the number of upper order statistics
where $\widehat{\zeta}_{i}$ denotes an estimator of the tail parameter $\eta$ computed from the $i$ largest order statistics. Throughout this work, we fix $\theta=0.3.\vskip 3.0pt plus 1.0pt minus 1.0pt$
This choice is supported by an extensive empirical calibration based on a preliminary study of $\theta$ over the interval $[0,$ $0.5],$ which systematically examined the sensitivity of the procedure to the tuning parameter and showed that $\theta=0.3$ consistently provides a favorable compromise in terms of bias and mean squared error across the considered scenarios. Smaller values of $\theta$ tend to oversmooth the variability among the upper order statistics, thereby masking relevant tail information, whereas larger values overemphasize extreme fluctuations, resulting in increased variance and reduced numerical stability.
4.6. Performance evaluation
Absolute bias (Bias) and mean squared error (MSE) are computed over a large number of Monte Carlo replications, providing a reliable assessment of the finite-sample behavior of the estimators.
Graphical representations are used to illustrate the combined effects of contamination, censoring, and the robustness parameter $\alpha$ on tail index estimation accuracy, with particular emphasis on stability with respect to the number of upper order statistics.
4.7. Simulation Results
This section is organized into two complementary parts. We first examine the behavior of the proposed estimator when contamination affects the latent data prior to the censoring mechanism. We then consider a more realistic setting in which contamination is introduced after censoring, when only censored observations are available.
This structure allows for a direct and transparent comparison between two fundamentally different contamination mechanisms.
4.7.1. Contamination introduced ”before” the censoring mechanism
We investigate the finite-sample performance of the proposed MDPD-based estimator under contamination introduced before the censoring mechanism. Figures 1–8 summarize the results for the Burr (Scenario S1) and Fréchet (Scenario S2) models across different tail indices, censoring levels, and contamination proportions.
Scenario S1 (Burr distribution)
Case $\gamma_{1}=0.3$ (Figures 9.19.2):
These figures report the bias and mean squared error (MSE) of the competing estimators for a Burr distribution with moderate tail heaviness $\left(\gamma_{1}=0.3\right),$ under two censoring levels, $p=0.55$ and $p=0.7.$
In the absence of contamination $\left(\epsilon=0\right),$ all estimators exhibit comparable performance for intermediate values of the number of upper order statistics $k.$ The classical estimators $\gamma_{1}^{(EFG)}$ and $\gamma_{1}^{(W)}$ slightly outperform the robust alternatives in terms of MSE, reflecting their higher efficiency under the correctly specified model. The MDPD estimator with a small robustness parameter $\left(\alpha=0.1\right)$ behaves very similarly, indicating that the loss of efficiency induced by the divergence remains limited in uncontaminated settings.
When contamination is introduced ($\epsilon=0.15$ and $\epsilon=0.40$), a substantial degradation of the classical estimators is observed, particularly for large values of $k$. Their bias increases rapidly, leading to a strong inflation of the MSE. In contrast, the MDPD estimators with $\alpha=0.3$ and $\alpha=0.5$ display remarkable stability, both in terms of bias and MSE, across a wide range of $k.$ This robustness is reflected by noticeably flatter curves, indicating a reduced sensitivity to the choice of the threshold parameter.
Increasing the proportion of uncensored observations ($p=0.7,$ Figure 9.2) improves the overall performance of all estimators. However, the qualitative conclusions remain unchanged: the MDPD estimators clearly dominate the classical approaches as soon as contamination becomes non-negligible.
Case $\gamma_{1}=0.5$ (Figures 9.39.4):
These figures corresponds to a heavier-tailed Burr distribution with $\gamma_{1}=0.5,$ which represents a more challenging estimation setting.
Even in the uncontaminated case, the classical estimators exhibit increased variability, especially for large values of $k.$ This instability is greatly amplified under contamination. For $\epsilon=0.40,$ both $\gamma_{1}^{(EFG)}$ and $\gamma_{1}^{(W)}$ become severely biased, resulting in very large MSE values and unreliable estimation.
By contrast, the MDPD estimators with $\alpha\geq 0.3$ remain stable over a broad range of $k,$ with controlled bias and substantially lower MSE. The protective effect of the robustness parameter $\alpha$ is particularly evident in this setting: larger values of $\alpha$ yield stronger resistance to contamination, at the cost of a moderate increase in variance for small $k.$ A higher censoring level $\left(p=0.55\right)$ exacerbates estimation difficulty, but does not alter the relative superiority of the robust procedures.
Scenario S2 (Fréchet distribution)
Case $\gamma_{1}=0.3$ (Figures 9.59.6):
These figures present the results for the Fréchet model with $\gamma_{1}=0.3.$ The conclusions are largely consistent with those obtained under the Burr model, highlighting the robustness of the proposed approach with respect to the underlying distribution.
In the absence of contamination, the differences between estimators remain moderate. However, as soon as contamination is introduced, the classical estimators become highly sensitive to extreme outliers, leading to rapidly increasing bias and MSE. In contrast, the MDPD estimators, particularly for $\alpha=0.3$ and $\alpha=0.5,$ exhibit strong resistance to contamination and clearly outperform the benchmark methods over the entire range of $k.\vskip 3.0pt plus 1.0pt minus 1.0pt$
Case $\gamma_{1}=0.5$ (Figures 9.79.8):
@@ 4.7.1. Contamination introduced ”before” the censoring mechanism
These figures illustrate the most adverse configuration, combining heavy tails, censoring, and severe contamination. In this setting, the classical estimators perform very poorly, with extremely unstable behavior and inflated MSE, even for moderate values of $k.$
The MDPD estimators remain reliable and display a markedly improved stability. In particular, the choice $\alpha=0.5$ provides the strongest protection against contamination, yielding nearly flat bias and MSE curves. Although this choice entails a slight loss of efficiency in ideal conditions, it appears especially suitable in practical situations where the presence of outliers in the upper tail cannot be ruled out.
Overall discussion (before case)
The simulation results consistently highlight the following points. First, in the absence of contamination, the MDPD estimator with a small robustness parameter remains competitive with classical estimators. Second, in contaminated settings, the classical methods deteriorate rapidly, whereas the MDPD estimators maintain controlled bias and MSE. Third, the tuning parameter $\alpha$ offers a transparent and effective way to balance efficiency and robustness, with $\alpha=0.3$ providing an excellent compromise in most scenarios. Finally, the MDPD estimators exhibit reduced sensitivity to the choice of the number of upper order statistics, which constitutes a major practical advantage.
Overall, these findings demonstrate that the proposed MDPD-based approach provides a robust and reliable alternative for tail index estimation under random censoring, particularly in realistic settings involving contaminated extremes.
4.7.2. Contamination introduced “after” the censoring mechanism
We now investigate the finite-sample performance of the proposed MDPD-based tail index estimator when contamination is introduced after the censoring mechanism. In contrast to the previous setting, this framework corresponds to a more adverse situation in which outliers directly affect the observed censored sample itself, thereby making the estimation problem intrinsically more challenging than in the case of contamination occurring before censoring.
Scenario S1 (Burr distribution)
Case $\gamma_{1}=0.3$ (Figures 9.99.10):
For moderately heavy tails, post-censoring contamination already induces a noticeable deterioration in performance. For both censoring levels $p=0.55$ and $p=0.7,$ the classical estimators $\gamma_{1}^{(EFG)}$ and $\gamma_{1}^{(W)}$ exhibit pronounced instability as soon as contamination is present. Their bias increases rapidly with $k$, resulting in inflated MSE curves, particularly under severe contamination $\left(\epsilon=0.40\right)$. This behavior reflects the high sensitivity of classical tail estimators to perturbations affecting the observed extremes.
The MDPD estimators display a more contrasted behavior than in the ”before” case. While a small robustness parameter $\left(\alpha=0.1\right)$ is no longer sufficient to fully protect against contamination, larger values ($\alpha=0.3$ and $\alpha=0.5$) still provide a clear stabilization effect. These robust versions yield flatter bias curves and reduced MSE over a broad range of $k,$ although the optimal stability region becomes narrower. This contraction of the stability region highlights the increased difficulty of the post-censoring contamination setting. Increasing the proportion of uncensored observations $(p=0.7)$ improves the overall performance, without altering the relative advantage of the robust estimators.
Case $\gamma_{1}=0.5$ (Figures 9.119.12):
For heavier tails, the impact of post-censoring contamination is further amplified. The classical estimators deteriorate rapidly, showing large bias and very unstable MSE curves even for moderate values of $k.$ This confirms their strong sensitivity to outliers that directly affect the observed tail, especially in the presence of heavy-tailed distributions.
The MDPD estimators offer improved robustness, but their effectiveness now strongly depends on the choice of $\alpha.$ While $\alpha=0.3$ still provides reasonable stability, $\alpha=0.5$ clearly emerges as the most reliable option in this setting, yielding the smallest MSE and the least sensitivity to $k,$ at the cost of a moderate loss of efficiency for small $k.$ This illustrates the bias–variance–robustness trade-off inherent to the choice of the tuning parameter.
Scenario S2 (Fréchet distribution)
Case $\gamma_{1}=0.3$ (Figures 9.139.14):
The results obtained under the Fréchet model are fully consistent with those observed for the Burr distribution. Post-censoring contamination has a severe impact on the classical estimators, whose bias and MSE increase rapidly with $k,$ especially under strong contamination. The degradation is systematic and persists across the entire range of $k.$ In contrast, the MDPD estimators remain noticeably more stable, particularly for $\alpha\geq 0.3,$ highlighting the benefit of robustness in heavy-tailed models.
Case $\gamma_{1}=0.5$ (Figures 9.159.16):
This configuration represents the most adverse scenario, combining heavy tails, censoring, and contamination introduced after censoring. In this case, the classical estimators perform very poorly, exhibiting extremely unstable behavior and highly inflated MSE values across almost the entire range of $k.$ Even small deviations from the uncontaminated model are sufficient to cause severe estimation failures.
The MDPD estimators clearly dominate in this setting. In particular, the choice $\alpha=0.5$ provides the strongest resistance to contamination, yielding relatively stable bias and MSE curves even under severe contamination. Although this choice entails a loss of efficiency in less contaminated regions, the gain in robustness is substantial and practically decisive. This confirms the ability of the MDPD approach to adapt to highly adverse contamination regimes through an appropriate tuning of $\alpha.$
Overall discussion (After case)
Overall, contamination introduced after the censoring mechanism proves to be significantly more harmful than contamination occurring before censoring. While all estimators are affected, the classical methods deteriorate rapidly and become unreliable in most contaminated scenarios. The MDPD estimators retain a clear advantage, but require larger values of the robustness parameter to maintain stability. These findings underscore the need for stronger robustness when contamination directly impacts the observed censored data, rather than the latent variable.
4.7.3. Global comparison: contamination before vs after censoring.
A clear contrast emerges when comparing contamination introduced before and after the censoring mechanism. When contamination occurs prior to censoring, its impact is partially mitigated by the censoring process, which tends to attenuate the influence of extreme outliers on the observed data. In this setting, the MDPD estimators with moderate robustness parameters already provide a substantial improvement over the classical approaches, while remaining competitive in uncontaminated cases.
In contrast, contamination introduced after censoring directly affects the observed extremes and leads to a much more severe degradation of estimation accuracy. This scenario proves to be considerably more challenging, as evidenced by the pronounced instability and rapidly increasing bias and MSE of the classical estimators across most configurations. Even moderate levels of contamination are sufficient to render the classical procedures unreliable, particularly for large values of the number of upper order statistics.
The MDPD estimators retain a clear advantage in both settings, but the comparison highlights an important difference in their behavior. While small robustness parameters are often sufficient when contamination occurs before censoring, stronger robustness (larger values of $\alpha$) becomes necessary when contamination is introduced after censoring. This clearly illustrates the critical role of the tuning parameter in adapting the estimator to both the timing and the severity of contamination.
Overall, these findings emphasize that post-censoring contamination constitutes a substantially more adverse scenario for tail index estimation. They also confirm that the proposed MDPD-based approach offers a flexible and reliable framework, capable of accommodating both situations through an appropriate choice of the robustness parameter, thereby providing a practical tool for inference under random censoring and contaminated extremes.
@@ 5.2. Estimation of the proportion $p$ of non-censored extremes
To justify the application of the MDPD estimator under weak censoring, we estimate the proportion of uncensored extreme observations $p.$ The adaptive algorithm of Reiss and Thomas is applied to select the optimal number of upper order statistics $k^{\ast}$ for estimating $p.$ Following empirical calibration, we set $\theta=0.3$, providing a robust compromise between stability and sensitivity to extreme values. Using this method, the optimal $k^{\ast}$ for estimating $p$ is $51,$ yielding an estimated $p=0.76,$ confirming that weak censoring $(p>1/2)$ holds in this dataset.
Note: In this real dataset, we do not apply artificial contamination, because the MDPD estimator is already robust both theoretically and as confirmed in simulations. Moreover, there is no clear evidence that the insurance data are contaminated, and introducing artificial contamination could misrepresent the true distribution. The dataset is treated as observed, reflecting realistic insurance losses, including naturally occurring extreme values.
@@ 5.3. Tail index estimation using MDPD
The same Reiss-Thomas algorithm is used to select $k^{\ast}$ for tail index estimation, resulting in $k^{\ast}=73,$ which is kept fixed across all contamination scenarios. The MDPD estimator is computed for four robustness parameter values:
$\alpha\in\{0.01,0.1,0.3,0.5\}.$
@@ 5.4. Contamination scenarios
    Remark: In contrast to simulations, contamination is not applied to real data, because:
@@ 5.6. Discussion and key takeaways
In conclusion, Section $5$ highlights that the MDPD approach can be applied directly to real datasets, capturing extreme events effectively while retaining robustness, efficiency, and interpretability.
@@ 5.7. Australian AIDS survival data (strong censoring)
The proportion of uncensored observations, denoted $p,$ is estimated using the Reiss–Thomas adaptive algorithm. The optimal number of upper order statistics for estimating $p$ is $k^{\ast}=162,$ yielding an estimated $p\approx 0.29.$ This indicates strong right censoring, with less than one-third of the observations being uncensored.
Under such strong censoring, the theoretical assumptions for the MDPD tail index estimator ($p>1/2$) are not satisfied. Consequently, the estimator’s performance may degrade, showing increased variability and potential bias. Attempts to compute the tail index using the Kaplan–Meier-based Worms–Worms estimator ($\gamma_{1}^{(W)}$) or the Nelson–Aalen-based MDPD estimator with $\alpha=0$ ($\gamma_{1}^{(MNS)}$) either fail to converge or produce unreliable results. Only the adapted Hill-type estimator ($\gamma_{1}^{(EFG)}$) provides estimates; however, it lacks robustness, inheriting the classical Hill estimator’s sensitivity to extreme observations.

### v6

@@ 4.4. Simulation scenarios and settings
Although infinitely many combinations of target, censoring, and contamination distributions could be considered, we restrict attention to two representative scenarios. These scenarios are standard in the extreme value literature and allow for a clear and informative assessment of both efficiency and robustness under censoring and contamination.
@@
In Scenario S1 (Burr/Burr/Burr), the target variable $X$ follows a Burr distribution, the censoring variable $C$ is also Burr-distributed, and contamination is introduced through a Burr distribution with parameters $\eta_{c}=0.25$ and $\gamma_{c}\in\{0.6,0.8\}.$
@@
In Scenario S2 (Fréchet/Fréchet/Fréchet), the target variable $X$ follows a Fréchet distribution, the censoring mechanism is Fréchet-distributed, and contamination is generated from a Fréchet distribution with tail index $\gamma_{c}\in\{0.6,0.8\}.$
These configurations allow us to examine robustness across different tail behaviors while maintaining a coherent and interpretable simulation framework. Contamination intensity is controlled by the parameter $\epsilon\in(0,1),$ which governs the proportion of observations affected by contamination in the upper tail.
For each scenario, contamination levels $\epsilon\in\{0,0.15,0.40\}$ are considered. These values correspond respectively to an uncontaminated setting, a moderately contaminated setting, and a severely contaminated configuration designed to stress-test robustness properties while remaining within a controlled and interpretable simulation framework.
4.5. Simulation results under pre-censoring contamination
This section investigates the finite-sample performance of the proposed MDPD-based tail index estimator under pre-censoring contamination, that is, when contamination affects the latent lifetime variable prior to the censoring mechanism. As emphasized in Section 4, this contamination framework is statistically coherent with both parametric and nonparametric censored models, preserves the independent censoring structure, and provides a meaningful benchmark for assessing robustness with respect to perturbations of the underlying tail distribution.
Scenario S1 (Burr distribution).
Case $\gamma_{1}=0.3$ (Figures 10.110.2). In the absence of contamination $\left(\epsilon=0\right),$ all estimators exhibit comparable behavior for intermediate values of the number of upper order statistics $k.$ As expected, classical estimators such as $\widehat{\gamma}_{1}^{(\mathrm{EFG})},$ $\widehat{\gamma}_{1}^{(\mathrm{W})},$ and $\widehat{\gamma}_{1}^{(\mathrm{MNS})}$ achieve good efficiency in this ideal setting, although their performance remains sensitive to the choice of the threshold parameter $k,$ a well-known feature of tail index estimation.
When contamination is introduced, a markedly different behavior emerges. Classical estimators deteriorate rapidly as $k$ increases, displaying pronounced bias and instability even for moderate contamination levels. This degradation reflects their intrinsic sensitivity to extreme observations generated by the heavier-tailed contaminating component. In contrast, the MDPD estimators exhibit a clear robustness effect: while small values of $\alpha$ retain some sensitivity to contamination, the estimators corresponding to $\alpha\geq 0.3$ remain stable over a wide range of $k,$ with substantially reduced bias and MSE. This clearly illustrates the role of the tuning parameter $\alpha$ in controlling the influence of outlying extremes and stabilizing inference.
Case $\gamma_{1}=0.5$ (Figures 10.310.4). For heavier tails, the adverse effect of contamination becomes more pronounced. Even in the uncontaminated case, classical estimators show increased variability, reflecting the intrinsic difficulty of tail index estimation in heavier-tailed regimes. Under contamination, these estimators become severely biased and unstable, with performance degrading rapidly as $k$ grows.
By contrast, the MDPD estimators maintain a remarkable degree of stability. In particular, estimators with $\alpha\geq 0.3$ exhibit nearly flat trajectories across a broad range of $k,$ indicating strong robustness with respect to both tail heaviness and contamination intensity. These results confirm that robustification becomes increasingly critical as the tail index increases, that is, when extremes play a dominant role.
Scenario S2 (Fréchet distribution).
Case $\gamma_{1}=0.3$ (Figures 10.510.6). The qualitative behavior observed in the Burr case carries over to the Fréchet setting. In the absence of contamination, all estimators perform satisfactorily for moderate values of $k,$ although classical estimators again display noticeable sensitivity to the choice of the threshold.
Under contamination, classical estimators experience a substantial loss of accuracy, with increasing bias and variability. This confirms that their lack of robustness is not distribution-specific but persists across different heavy-tailed models. In contrast, the MDPD estimators demonstrate a robust behavior similar to that observed in Scenario S1, with larger values of $\alpha$ yielding improved stability and resistance to contamination-induced distortions.
@@ 4.5. Simulation results under pre-censoring contamination
Case $\gamma_{1}=0.5$ (Figures 10.710.8). This configuration represents the most adverse scenario, combining heavy tails with severe contamination. In this setting, classical estimators perform particularly poorly, often exhibiting dramatic bias and erratic behavior across the range of $k$ values.
The MDPD estimators, however, remain reliable even in this challenging regime. Although a moderate increase in variance is observed for very large values of $\alpha,$ this effect is largely compensated by a substantial reduction in bias, so that the overall MSE remains significantly lower than that of the classical competitors. This highlights the practical advantage of the MDPD approach in extreme scenarios where standard methods effectively break down.
Overall discussion.
Overall, the simulation results provide strong empirical evidence that the proposed MDPD-based tail index estimator offers a robust and reliable alternative for censored extreme value analysis. While classical estimators may retain efficiency under ideal conditions, their performance deteriorates sharply in the presence of contaminated extremes. In contrast, the MDPD estimators achieve a favorable efficiency–robustness trade-off, particularly for moderate to large values of the tuning parameter $\alpha.$
Taken together, these findings demonstrate that the proposed methodology is well suited for realistic applications, where both censoring and contamination are unavoidable and robustness with respect to extreme observations is a key requirement.
@@ 5.1. Insurance loss data (weak censoring)
Before proceeding to tail index estimation, it is essential to assess the effective amount of tail information available in the presence of censoring. In particular, we first estimate the proportion of uncensored extreme observations in order to verify that the weak censoring assumption required by the proposed estimator is satisfied.
@@ 5.2. Estimation of the proportion $p$ of non-censored extremes
To justify the application of the MDPD estimator under weak censoring, we estimate the proportion $p$ of uncensored extreme observations. This quantity plays a central role, as it determines whether a sufficient number of extreme observations remain fully observed to support reliable tail inference.
The adaptive algorithm of Reiss and Thomas is applied to select the optimal number of upper order statistics $k^{\ast}$ used for estimating $p$. This data-driven procedure allows us to balance stability and sensitivity in the tail region, while accounting for the presence of censoring.
5.3. Selection of the number of upper order statistics
where $\widehat{\zeta}_{i}$ denotes an estimator of the tail parameter $\eta$ computed from the $i$ largest order statistics.
Throughout this work, we fix $\theta=0.3.$ This choice is supported by an extensive empirical calibration based on a preliminary study of $\theta$ over the interval $[0,$ $0.5],$ which systematically examined the sensitivity of the procedure to the tuning parameter and showed that $\theta=0.3$ consistently provides a favorable compromise in terms of bias and mean squared error across the considered scenarios. Smaller values of $\theta$ tend to oversmooth the variability among the upper order statistics, thereby masking relevant tail information, whereas larger values overemphasize extreme fluctuations, resulting in increased variance and reduced numerical stability.
Applying this procedure to the insurance loss data yields an optimal value $k^{\ast}=51$ for estimating $p.$ The resulting estimate $\widehat{p}=0.76,$ confirms that the weak censoring condition $(p>1/2)$ holds in this dataset, validating the applicability of the proposed MDPD-based approach.
Note. In this real-data analysis, we do not introduce artificial contamination. The MDPD estimator is already robust both from a theoretical perspective and as confirmed by the simulation results. Moreover, there is no clear empirical evidence suggesting contamination in the insurance data, and artificially modifying the observations could distort the underlying loss distribution. The dataset is therefore analyzed as observed, reflecting realistic insurance losses, including naturally occurring extreme values.
@@ 5.4. Tail index estimation using MDPD
Having verified the weak censoring assumption and determined an appropriate threshold, we now proceed to tail index estimation. The same Reiss–Thomas algorithm is employed to select the number of upper order statistics for tail index estimation, leading to $k^{\ast}=73,$ This value is kept fixed across all robustness settings to facilitate comparison.
The MDPD estimator is computed for four robustness parameter values:
$\alpha\in\{0.01,0.1,0.3,0.5\},$
allowing us to assess the impact of robustness on tail index estimation in a realistic insurance context.
@@ 5.5. Contamination scenarios
    Note. In contrast to simulations, contamination is not applied to real data, because:
@@ 5.7. Discussion and key takeaways
In conclusion, Section 5 highlights that the MDPD approach can be applied directly to real datasets, capturing extreme events effectively while retaining robustness, efficiency, and interpretability.
@@ 5.8. Australian AIDS survival data (strong censoring)
The proportion of uncensored observations, denoted $p,$ is estimated using the Reiss–Thomas adaptive algorithm. The optimal number of upper order statistics for estimating $p$ is $k^{\ast}=162,$ yielding an estimated $\widehat{p}=0.29.$ This indicates strong right censoring, with less than one-third of the observations being uncensored.
Under such strong censoring, the theoretical assumptions for the MDPD tail index estimator $(p>1/2)$ are not satisfied. Consequently, the estimator’s performance may degrade, showing increased variability and potential bias. Attempts to compute the tail index using the Kaplan–Meier-based Worms–Worms estimator ($\widehat{\gamma}_{1}^{(\mathrm{W})}$) or the Nelson–Aalen-based MDPD estimator with $\alpha=0$ ($\gamma_{1}^{(\mathrm{MNS})}$) either fail to converge or produce unreliable results. Only the adapted Hill-type estimator ($\widehat{\gamma}_{1}^{(\mathrm{EFG})}$) provides estimates; however, it lacks robustness, inheriting the classical Hill estimator’s sensitivity to extreme observations.